It is determined that a force F of 201 N at an ang
Answer #1

Normal force exerted by the wall on the block

N = F*cos(10.7)

= 201*cos(10.7)

= 197.5 N

now apply, Fnety = 0

fs_max - m*g = 0

mue_s*N - m*g = 0

mue_s = m*g/N

= 11.2*9.8/197.5

= 0.5557 <<<<<<<<----------------Answer